Solve each of the following systems by substitution.
5x-2y=-5
y-5x=3

Answer by tutorcecilia(2152)

5x-2y=-5
y-5x=3
Substitute one value for the other by solving for either x or y.
.
y-5x=3
y=5x+3
.
5x-2y=-5
5x-2(5x+3)=-5
5x-10x-6=-5
-5x-6=-5
-5x=-5+6
-5x=1
x=-1/5 [Plug it back into the other equation]
.
y-5(-1/5)=3
y+1=3
y=2
.
Plug the values of (x=(-1/5) and (y=2) into either equation:
5x-2y=-5
5(-1/5)-2(2)=-5
-5=-5 [checks out]